OpenStudy (anonymous):

Calc (derivatives) the question is y=x-6-(5/x)-(x/6)+e^(5x) I got 3+e^(5x), but I don't think that is right. I took the derivative of (-5/x)^(1), which would be 1, but I am kind of lost as to what to do. Could I instead do -5x^(-1)?

OpenStudy (anonymous):

no it is not right

OpenStudy (anonymous):

does it really say \[\large y=x-6-\frac{5}{x}-\frac{x}{6}+e^{5x}\]?

OpenStudy (anonymous):

like this ? \(\Huge y=x-6-\frac{5}{x}-\frac{x}{6}+e^{5x}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( \frac{ u }{v } \right)=\frac{ vu'-uv' }{v^2 }\]

OpenStudy (anonymous):

What I am doing now, after restarting, is (just for the fractions) -5(x^(-1), which means i have to use the multplication rule

OpenStudy (anonymous):

ok first of all \[x-\frac{x}{6}=\frac{5}{6}x\]

OpenStudy (anonymous):

and the derivative of \(\frac{5}{6}x\) is \(\frac{5}{6}\)

OpenStudy (anonymous):

or\[\frac{ d }{ dx }\left( \frac{ 1 }{x^n } \right)=\frac{ d }{dx }\left( x ^{-n} \right)=-nx ^{-n-1}\]

OpenStudy (anonymous):

the derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\) so the derivative of \(-\frac{5}{x}\) is \(\frac{5}{x^2}\)

OpenStudy (anonymous):

use no rule use your memory the derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\) and by the chain rule, the derivative of \(\frac{1}{f(x)}\) is \(-\frac{f'(x)}{f^2(x)}\)

OpenStudy (anonymous):

My teacher wants us to show our work. How did you get -5/x =5/(x^(2)?

OpenStudy (anonymous):

explanation above if showing work is really necessary, either you have to show \[\lim_{h\to 0}\frac{\frac{-5}{x+h}+\frac{5}{x}}{h}=\frac{5}{x^2}\] which is the only work to do if you get to use shortcut rules, then the quickest shortcut is to remember it the reciprocal is a very common function, and the derivative never changes

OpenStudy (anonymous):

wait, I think I have it. is the answer:\[-5+\frac{ 5 }{ x^2 }+\frac{ x }{ 6^2 } +e ^{5x}\]

OpenStudy (anonymous):

there are two you should simply remember without using any rules one is the rule for square roots \[\frac{d}{dx}\sqrt{x}]=\frac{1}{2\sqrt{x}}\] and by the chain rule \[\frac{d}{dx}[\sqrt{f(x)}]=\frac{f'(x)}{2\sqrt{f(x)}}\] those should be immediate

OpenStudy (anonymous):

no combine like terms first

OpenStudy (anonymous):

\[-\frac{x}{6}=-\frac{1}{6}x\]

OpenStudy (anonymous):

and so the derivative of \[-\frac{1}{6}x\] is \[-\frac{1}{6}\]

OpenStudy (anonymous):

also, by the chain rule, the derivative of \(e^{5x}\) is \[e^{5x}\frac{d}{dx}[5x]\] i.e. \[5e^{5x}\]

OpenStudy (anonymous):

oh, I see

OpenStudy (anonymous):

I thought that the derivative of e^(5x) was just e^(5x)

OpenStudy (anonymous):

final answer should be \[\frac{5}{6}+\frac{5}{x^2}+5e^{5x}\] i think the \[x-\frac{x}{6}\] was a trick to make sure you were paying attention i guess most times you do not see a function without the like terms combined

OpenStudy (anonymous):

chain rule for these, and it is easy to forget when you have things like \[\sin(2x)\] for example the derivative of \[\sin(2x)\] is \[2\cos(2x)\] don't forget the 2 out front

OpenStudy (anonymous):

Ok, I think i am getting it.

OpenStudy (anonymous):

I got it, thanks!!!